# SparseVLM: Visual Token Sparsification for Efficient Vision-Language Model Inference

### v3

@@ Abstract
In vision-language models (VLMs), visual tokens usually bear a significant amount of computational overhead despite sparsity of information in them when compared to text tokens. To address this, most existing methods learn a network to prune redundant visual tokens using certain training data. Differently, we propose a text-guided training-free token optimization mechanism dubbed SparseVLM that eliminates the need of extra parameters or fine-tuning costs. Given that visual tokens complement text tokens in VLM’s linguistic reasoning, we select relevant text tokens to rate the significance of visual tokens using self-attention matrices and, then, prune visual tokens using the proposed strategy to maximize sparsity while retaining information. In particular, we introduce a rank-based strategy to adaptively determine the sparsification ratio for each layer, alongside a token recycling method that compresses pruned tokens into more compact representations. Experimental results show that SparseVLM increases the efficiency of various VLMs in a number of image and video understanding tasks. For example, LLaVA when equipped with SparseVLM achieves 54% reduction in FLOPs, 37% decrease in CUDA latency while maintaining 97% of its original accuracy. Our code is available at https://github.com/Gumpest/SparseVLMs.
@@ 1 Introduction
Figure 1: Comparison of visual token sparsification methods. Unlike previous methods with text-agnostic visual sparsification (c) e.g., VocoLLaMA (Ye et al., 2024), our SparseVLM (b) is guided by question prompts to select relevant visual patches from the image (a).
Benefiting from advancements in large language models (LLMs) (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Peng et al., 2023; Bi et al., 2024), the realm of vision-language models (VLMs) has undergone significant progress. To combine visual signals with textual semantics, the mainstream practice in VLMs (Team et al., 2023; Bai et al., 2023; Chen et al., 2023; Li et al., 2024c, 2023a) employs sequential visual representation, where images are extracted into visual tokens and sent into an LLM decoder. With modal alignment and instruction fine-tuning (Du et al., 2021; Liu et al., 2023a; Zhu et al., 2023b), recent VLMs successfully adapt LLMs to the vision domain and inherit their perception and reasoning abilities.
Despite the promising performance, further incorporation of visual tokens inevitably introduces a huge memory and computational overhead when compared to LLMs, particularly for high-resolution images (Li et al., 2024c) and long videos (Lin et al., 2023). For instance, a $672\times 672$ image in LLaVA (Liu et al., 2024) yields $2304$ visual tokens that span over half of the context length. However, the information in images is typically more sparse than in natural languages (Marr, 2010), resulting in inefficiency when naïvely processing both modalities. To address this, existing methods extract more compact image representations by modifying the image encoder or projector (Alayrac et al., 2022; Li et al., 2024b; Dai et al., 2023; Cha et al., 2024). While some recent works further sparsify visual tokens during the decoding (Ye et al., 2024; Chen et al., 2024; Cai et al., 2024), they still ignore the guidance from the language tokens, which contradicts the multimodality paradigm. We argue that visual tokens should be sparsified adaptively based on the question prompt, as the model might focus on different parts (e.g., foreground or background) when dealing with various questions as shown in Figure 1. Furthermore, current approaches generally train a network to prune redundant visual tokens and require additional training data (Li et al., 2024b; Ye et al., 2024).
@@
The proposed method is simple yet practical. It can act as a plug-and-play module to improve the efficiency of VLMs without additional fine-tuning. Extensive experiments demonstrate that our SparseVLM effectively reduces computational overhead of various VLMs without sacrificing their performance in a wide range of image and video understanding tasks. For instance, LLaVA (Liu et al., 2024) when armed with SparseVLM achieves a 4.5×4.5$\times$4.5 × compression rate while maintaining $97\%$ of its original performance. Alternatively, the CUDA latency can decrease by $37\%$ with only a $0.9\%$ drop in accuracy. To investigate the effectiveness of our method in video tasks, we further apply SparseVLM to VideoLLaVA (Lin et al., 2023) to compress frames with temporal dimension. Without complex design changes, SparseVLM can sparsify video frames into an adaptive number of visual tokens and outperform existing methods in video question-answering benchmarks. Our approach consistently outperforms prior state-of-the-art FastV method (Chen et al., 2024) by $11.2-17.3\%$ on LLaVA, $9.2-20.4\%$ on MiniGemini, and $34.4\%$ on VideoLLaVA when both have similar latencies.
@@ 2 Related Work
Vision-Language Models. Recent works on vision-language models (Liu et al., 2023a; Chen et al., 2023; Li et al., 2024c) improve multimodal comprehension and generation by processing longer visual token sequences. Moreover, the usage of higher-resolution images inevitably entails an exponential growth in the length of visual sequences. For example, LLaVA typically encodes $336\times 336$ images into $576$ tokens (Liu et al., 2024) with up to $672\times 672$ maximum resolution using $2880$ token sequences (Liu et al., 2023a). Similarly, mini-Gemini-HD (Li et al., 2024c) converts $1536\times 1536$ high resolution and $672\times 672$ low resolution images into $2880$ visual tokens. Moreover, comprehending videos or multiple images leads to increased token allocations for visual signals. For instance, the VideoLLaVA (Lin et al., 2023) and VideoPoet (Kondratyuk et al., 2023) use thousands of tokens to encode multiple image frames. However, large number of visual tokens results in a computational bottleneck. Further research on sparsification is urged to further unleash VLM capabilities.
Visual Compression for VLMs. Compression of visual tokens is necessary because, on the one hand, their quantity is usually tens to hundreds of times that of language tokens. On the other hand, visual signals are inherently more sparse in information when compared to texts that have been produced by humans (Marr, 2010). Past efforts to address the above problem can be categorized into two directions. The first one centers on the compression of a vision tower or an efficient projection of vision modality. For instance, LLaMA-VID (Li et al., 2024b) exploits the Q-Former with the context token while DeCo (Yao et al., 2024) employs an adaptive pooling to downsample the visual tokens at the patch level. Methods that belong to the second direction (Ye et al., 2024; Chen et al., 2024; Cai et al., 2024) go deeper into the text modality and sparsify visual tokens during the LLM decoding stage, but they still lack the guidance from the text tokens. In this paper, SparseVLM takes note of this limitation and improves performance upon it.
@@ 3.1 Preliminary: Attention in VLM Decoders
Often, the matrix ${\bm{A}}$ cannot be directly accessed due to FlashAttention-type (Dao et al., 2022) optimizations. Therefore, we develop an approach to extract ${\bm{A}}$ while maintaining compatibility with the FlashAttention when applying our sparsification. Please refer to the Appendix A.2.
@@ 3.4 Theoretical Analysis of Computational Complexity
We consider the computation of multi-head attention and feed-forward network (FFN) modules in the FLOPs estimation. Assuming $N$ is the number of pruned tokens, $D$ is the hidden state size, which is the same as the intermediate size in FFN, the FLOPs for one Transformer layer can be reduced by $6(N-C)D^{2}+2(N-C)^{2}D$. Besides, our partial step introduces minimal computation with the details provided in Appendix A.3. Thus, we estimate the FLOPs savings as the reduction part minus the additional overhead:
@@ 4.1 Image Understanding Tasks
Implementation Details. We verify SparseVLM on three VLM frameworks: LLaVA (Liu et al., 2024), Mini-Gemini (MGM) (Li et al., 2024c), and Qwen2-VL (Bai et al., 2023). LLaVA-1.5 employs CLIP-pretrained ViT-L as the visual tower, MGM further introduces a LAION-pretrained ConvNeXt-L (Liu et al., 2022) for high-resolution refinement, while Qwen2-VL owns dynamic resolution encoder.
Main Results. In Table 1, we present the performance of SparseLLaVA (LLaVA equipped with SparseVLM) on image understanding benchmarks. To intuitively assess the performance, we provide the results by percentage format for comparative analysis, and the accuracy of the vanilla model with the 100% upper limit. We set $3$ vision token count configurations ($192$, $128$, and $64$) to check the advantages of SparseVLM comprehensively. When pruning from $576$ to $192$ tokens, the SparseLLaVA only decreases the average accuracy by $0.9\%$ without additional training and exceeds ToMe (Bolya et al., 2022) $10.2\%$. When only $64$ tokens are kept, our method outperforms FastV (Chen et al., 2024) by a significant margin of 17.3%, while ToMe performs worst due to its direct merging. Furthermore, we also compare the recent method PDrop (Xing et al., 2024) training-free version, which has lower FLOPs computation. However, our method outperforms it in accuracy and latency, which are the most crucial metrics for practical deployment.
@@ 4.2 Video Understanding Tasks
Table 3: The results of Video-LLaVA with SparseVLM on video question answering task. The original number of video tokens is $2048$, while our experiment collectively prunes it down to $198$ tokens. FastV (Chen et al., 2024) is included for comparison.
@@
Video-LLaVA 47.1 69.8 56.7 43.1 100.0%
100% 100% 100% 100%
FastV 23.1 38.0 19.3 30.6 52.1%
49.0% 54.4% 34.0% 71.0%
Ours 44.7 68.2 31.0 42.6 86.5%
94.9% 97.7% 54.7% 98.8%
Datasets. We test on four common video question answering benchmarks, TGIF-QA (Jang et al., 2017), MSVD-QA (Xu et al., 2017), MSRVTT-QA (Xu et al., 2017), and ActivityNet-QA (Yu et al., 2019), where video-question pairs are massively disproportional in length. We use the evaluation tool of Video-ChatGPT (Maaz et al., 2023).
Implementation Details. We directly apply our SparseVLM for Video-LLaVA (Lin et al., 2023), which is composed of several key components, including language bind encoder $f^{v}_{M}$(Zhu et al., 2023a) for extracting features from raw visual inputs (e.g., images or videos), a language decoder model $f_{L}$ such as Vicuna (Touvron et al., 2023), a visual projection layer $f_{P}$, and a word embedding layer $f_{T}$.
Main Results. In Table 3, we set the Video-LLaVA with $2048$ video tokens as our upper bound for an overall average accuracy of $100.0\%$ and a score of $+0.00$. To make a fair comparison, we both preserve $198$ vision tokens ($90.3\%$ pruning ratio) for FastV (Chen et al., 2024) and SparseVLM. It is clear that our approach consistently outperforms FastV across all benchmarks, both in accuracy (Acc.) and GPT evaluation score. SparseVideoLLaVA achieves a total average accuracy of $86.5\%$, a significant 34.4% higher than $52.1\%$ of FastV. (From the GPT score perspective, SparseVLM only loses $0.17$ points compared to $1.02$ points of FastV.) These improvements suggest that when handling video modality containing temporal features, SparseVLM continues to deliver strong performance, generating accurate responses to diverse questions while utilizing significantly fewer tokens. This achieves an effective trade-off between inference efficiency and model performance.
@@ 5.1 Relevant Text Token Selection
We propose a selection mechanism to localize visually irrelevant text tokens to limit their negative effects in rating the significance of vision tokens. Here we conduct experiments to analyze the effects of the mechanism in Figure 5. Under the same number of vision tokens ($64$), we have 3 settings (using all tokens, only text tokens, and only text raters we select) with LLaVA (Liu et al., 2023a) to judge vision token candidates. In TextVQA (Singh et al., 2019), by building upon the text-aware manner, our mechanism improves the baseline (all tokens) by $0.8\%$, which validates that our extra selection is effective. Besides, we further outperform the vanilla text-aware method (only text tokens) by $\textbf{2.7}\%$ on POPE (Li et al., 2023b). The huge margin means POPE sparsification is quite sensitive to question prompts, and text guidance is necessary. In summary, text rater selection is general and improves the performance across scenarios.
@@ 5.2 Recycling of Pruned Tokens
To validate the effectiveness of our token recycling strategy, we perform ablation experiments on the LLaVA model (Liu et al., 2023a). The results are presented in Table 4. Across multiple sparsity ratios ($64$, $96$, $128$, $192$), our algorithm achieves a significant average performance improvement of $\textbf{1.2}\%$ and $\textbf{7.2}\%$ on TextVQA (Singh et al., 2019) and POPE (Li et al., 2023b), respectively. Notably, as the number of pruned vision tokens increases, the benefit brought by our recycling method increases. For instance, when pruning from $192$ to $64$ tokens, the pruned token recycling significantly boosts the accuracy from 1.5% to 17.7% on POPE. We argue that when the size of the deleted pool grows, the amount of lost information increases. Our method effectively recycles the lost information and compresses it into few slots using the proposed reconstruction mechanism.
@@ 5.3 Computational Efficiency
SparseVLM affords significant efficiency and storage gains for the inference process. We conduct a comparative analysis of CUDA time, and FLOPs on LLaVA-7B, and compare our method with the baseline method and FastV (Chen et al., 2024). As displayed in Table 1, we conduct an inference efficiency analysis on a single NVIDIA A100-80GB with identical lengths of text prompts and single-image inputs. Compared to the baseline model, SparseVLM achieves a significant reduction of $43.1\%$ in CUDA time and $62.8\%$ in FLOPs while keeping $96.7\%$ accuracy. Despite SparseVLM has a minimal overhead to calculate text raters and cluster-pruned vision tokens, it leads to fewer than FastV tokens with comparable accuracy. Additionally, SparseVLM saves $67\%$ cache memory compared to vanilla LLaVA (where 302.4MB is reduced to 100.8MB), while keeping $99.1\%$ accuracy. More efficiency visualization (e.g., efficiency on VideoLLaVA) can be found in the Appendix A.8.

### v4

@@ Abstract
In vision-language models (VLMs), visual tokens usually bear a significant amount of computational overhead despite sparsity of information in them when compared to text tokens. To address this, most existing methods learn a network to prune redundant visual tokens using certain training data. Differently, we propose a text-guided training-free token optimization mechanism dubbed SparseVLM that eliminates the need of extra parameters or fine-tuning costs. Given that visual tokens complement text tokens in VLM’s linguistic reasoning, we select relevant text tokens to rate the significance of visual tokens using self-attention matrices and, then, prune visual tokens using the proposed strategy to maximize sparsity while retaining information. In particular, we introduce a rank-based strategy to adaptively determine the sparsification ratio for each layer, alongside a token recycling method that compresses pruned tokens into more compact representations. Experimental results show that SparseVLM increases the efficiency of various VLMs in a number of image and video understanding tasks. Our code is available at https://github.com/Gumpest/SparseVLMs.
@@ 1 Introduction
Figure 1: Comparison of visual token sparsification methods. Unlike previous methods with text-agnostic visual sparsification (c) e.g., VocoLLaMA (Ye et al., 2025), our SparseVLM (b) is guided by question prompts to select relevant visual patches from the image (a).
Benefiting from advancements in large language models (LLMs) (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Peng et al., 2023; Zhang et al., 2024a), the realm of vision-language models (VLMs) has undergone significant progress. To combine visual signals with textual semantics, the mainstream practice in VLMs (Team et al., 2023; Bai et al., 2023; Chen et al., 2024b; Li et al., 2024c, 2023a) employs sequential visual representation, where images are extracted into visual tokens and sent into an LLM decoder. With modal alignment and instruction fine-tuning (Du et al., 2022; Liu et al., 2024a; Zhu et al., 2024b), recent VLMs successfully adapt LLMs to the vision domain and inherit their perception and reasoning abilities.
Despite the promising performance, further incorporation of visual tokens inevitably introduces a huge memory and computational overhead when compared to LLMs, particularly for high-resolution images (Li et al., 2024c) and long videos (Lin et al., 2024). For instance, a $672\times 672$ image in LLaVA (Liu et al., 2024b) yields $2304$ visual tokens that span over half of the context length. However, the information in images is typically more sparse than in natural languages (Marr, 2010), resulting in inefficiency when naïvely processing both modalities. To address this, existing methods extract more compact image representations by modifying the image encoder or projector (Alayrac et al., 2022; Li et al., 2024b; Dai et al., 2023; Cha et al., 2024). While some recent works further sparsify visual tokens during the decoding (Ye et al., 2025; Chen et al., 2024a; Shang et al., 2024), they still ignore the guidance from the language tokens, which contradicts the multimodality paradigm. We argue that visual tokens should be sparsified adaptively based on the question prompt, as the model might focus on different parts (e.g., foreground or background) when dealing with various questions as shown in Figure 1. Furthermore, current approaches generally train a network to prune redundant visual tokens and require additional training data (Li et al., 2024b; Ye et al., 2025; Cai et al., 2025).
@@
The proposed method is simple yet practical. It can act as a plug-and-play module to improve the efficiency of VLMs without additional fine-tuning. Extensive experiments demonstrate that our SparseVLM effectively reduces computational overhead of various VLMs without sacrificing their performance in a wide range of image and video understanding tasks. For instance, LLaVA (Liu et al., 2024b) when armed with SparseVLM achieves a 4.5×4.5$\times$4.5 × compression rate while maintaining $97\%$ of its original performance. Alternatively, the CUDA latency can decrease by $37\%$ with only a $0.9\%$ drop in accuracy. To investigate the effectiveness of our method in video tasks, we further apply SparseVLM to VideoLLaVA (Lin et al., 2024) to compress frames with temporal dimension. Without complex design changes, SparseVLM can sparsify video frames into an adaptive number of visual tokens and outperform existing methods in video question-answering benchmarks. Our approach consistently outperforms prior state-of-the-art FastV method (Chen et al., 2024a) by $11.2-17.3\%$ on LLaVA, $9.2-20.4\%$ on MiniGemini, and $14.7\%$ on VideoLLaVA when both have similar latencies.
@@ 2 Related Work
Vision-Language Models. Recent works on vision-language models (Liu et al., 2024a; Chen et al., 2024b; Li et al., 2024c) improve multimodal comprehension and generation by processing longer visual token sequences. Moreover, the usage of higher-resolution images inevitably entails an exponential growth in the length of visual sequences. For example, LLaVA typically encodes $336\times 336$ images into $576$ tokens (Liu et al., 2024b) with up to $672\times 672$ maximum resolution using $2880$ token sequences (Liu et al., 2024a). Similarly, mini-Gemini-HD (Li et al., 2024c) converts $1536\times 1536$ high resolution and $672\times 672$ low resolution images into $2880$ visual tokens. Moreover, comprehending videos or multiple images leads to increased token allocations for visual signals. For instance, the VideoLLaVA (Lin et al., 2024) and VideoPoet (Kondratyuk et al., 2024) use thousands of tokens to encode multiple image frames. However, large number of visual tokens results in a computational bottleneck. Further research on sparsification is urged to further unleash VLM capabilities.
Visual Compression for VLMs. Compression of visual tokens is necessary because, on the one hand, their quantity is usually tens to hundreds of times that of language tokens. On the other hand, visual signals are inherently more sparse in information when compared to texts that have been produced by humans (Marr, 2010). Past efforts to address the above problem can be categorized into two directions. The first one centers on the compression of a vision tower or an efficient projection of vision modality. For instance, LLaMA-VID (Li et al., 2024b) exploits the Q-Former with the context token while DeCo (Yao et al., 2024) employs an adaptive pooling to downsample the visual tokens at the patch level. Methods that belong to the second direction (Ye et al., 2025; Chen et al., 2024a; Wu et al., 2024) go deeper into the text modality and sparsify visual tokens during the LLM decoding stage, but they still lack guidance from the text tokens. In this paper, SparseVLM takes note of this limitation and improves performance upon it.
@@ 3.1 Preliminary: Attention in VLM Decoders
Often, the matrix ${\bm{A}}$ cannot be directly accessed due to FlashAttention-type (Dao et al., 2022) optimizations. Therefore, we develop an approach to extract ${\bm{A}}$ while maintaining compatibility with the FlashAttention when applying our sparsification. Please refer to the Appendix B.
@@ 3.4 Theoretical Analysis of Computational Complexity
We consider the computation of multi-head attention and feed-forward network (FFN) modules in the FLOPs estimation. Assuming $N$ is the number of pruned tokens, $D$ is the hidden state size, which is the same as the intermediate size in FFN, the FLOPs for one Transformer layer can be reduced by $6(N-C)D^{2}+2(N-C)^{2}D$. Besides, our partial step introduces minimal computation with the details provided in Appendix C. Thus, we estimate the FLOPs savings as the reduction part minus the additional overhead:
@@ 4.1 Image Understanding Tasks
Implementation Details. We verify SparseVLM on three VLM frameworks: LLaVA (Liu et al., 2024b), Mini-Gemini (MGM) (Li et al., 2024c), and Qwen2-VL (Bai et al., 2023). LLaVA-1.5 employs CLIP-pretrained ViT-L as the visual tower, MGM further introduces a LAION-pretrained ConvNeXt-L (Liu et al., 2022) for high-resolution refinement, while Qwen2-VL owns dynamic resolution encoder.
Main Results. In Table 1, we present the performance of SparseLLaVA (LLaVA equipped with SparseVLM) on image understanding benchmarks. To intuitively assess the performance, we provide the results by percentage format for comparative analysis, and the accuracy of the vanilla model with the 100% upper limit. We set $3$ vision token count configurations ($192$, $128$, and $64$) to check the advantages of SparseVLM comprehensively. When pruning from $576$ to $192$ tokens, the SparseLLaVA only decreases the average accuracy by $0.9\%$ without additional training and exceeds ToMe (Bolya et al., 2023) $10.2\%$. When only $64$ tokens are kept, our method outperforms FastV (Chen et al., 2024a) by a significant margin of 17.3%, while ToMe performs worst due to its direct merging. Furthermore, we also compare the recent method PDrop (Xing et al., 2025) training-free version, which has lower FLOPs computation. However, our method outperforms it in accuracy and latency, which are the most crucial metrics for practical deployment.
@@ 4.2 Video Understanding Tasks
Table 3: The results of Video-LLaVA with SparseVLM on video question answering task. The original number of video tokens is $2048$, while our experiment collectively prunes it down to $194$ tokens. FastV (Chen et al., 2024a) is included for comparison.
@@
Acc. Score Acc. Score Acc. Score Acc. Score Acc. Score
Video-LLaVA 18.9 2.54 72.0 3.95 57.1 3.45 43.6 3.81 47.9 3.44
100% +0.00 100% +0.00 100% +0.00 100% +0.00
FastV (ECCV24) 10.2 2.29 58.3 3.62 52.3 3.42 41.3 3.76 40.5 3.27
54.0% -0.34 81.0% -0.33 91.6% -0.03 94.7% -0.12
Ours 14.9 2.41 71.7 3.94 56.1 3.43 45.1 3.81 47.0 3.40
78.8% -0.13 99.6% -0.01 98.3% -0.02 103.4% -0.00
Datasets. We test on four common video question answering benchmarks, TGIF-QA (Jang et al., 2017), MSVD-QA (Xu et al., 2017), MSRVTT-QA (Xu et al., 2017), and ActivityNet-QA (Yu et al., 2019). Specifically, following FastV’s (Chen et al., 2024a) setup, we use the first 1000 samples per benchmark and score them using the Video-ChatGPT (Maaz et al., 2024) evaluation tool, acknowledging the characteristic length imbalances in these datasets.
Implementation Details. We directly apply our SparseVLM for Video-LLaVA (Lin et al., 2024), which is composed of several key components, including language bind encoder $f^{v}_{M}$(Zhu et al., 2024a) for extracting features from raw visual inputs (e.g., images or videos), a language decoder model $f_{L}$ such as Vicuna (Touvron et al., 2023), a visual projection layer $f_{P}$, and a word embedding layer $f_{T}$.
Main Results. In Table 3, we set the Video-LLaVA with $2048$ video tokens as our upper bound for an overall average accuracy of $100.0\%$ and a score of $+0.00$. To make a fair comparison, we both preserve $194$ vision tokens ($90.5\%$ pruning ratio) for FastV (Chen et al., 2024a) and SparseVLM. It is clear that our approach consistently outperforms FastV across all benchmarks, both in accuracy (Acc.) and GPT evaluation score. SparseVideoLLaVA achieves a total average accuracy of $95.0\%$, a significant 14.7% higher than $80.3\%$ of FastV. (From the GPT score perspective, SparseVLM only loses $0.04$ points compared to $0.17$ points of FastV.) These improvements suggest that when handling video modality containing temporal features, SparseVLM continues to deliver strong performance, generating accurate responses to diverse questions while utilizing significantly fewer tokens. This achieves an effective trade-off between inference efficiency and model performance.
@@ 5.1 Relevant Text Token Selection
We propose a selection mechanism to localize visually irrelevant text tokens to limit their negative effects in rating the significance of vision tokens. Here we conduct experiments to analyze the effects of the mechanism in Figure 5. Under the same number of vision tokens ($64$), we have 3 settings (using all tokens, only text tokens, and only text raters we select) with LLaVA (Liu et al., 2024a) to judge vision token candidates. In TextVQA (Singh et al., 2019), by building upon the text-aware manner, our mechanism improves the baseline (all tokens) by $0.8\%$, which validates that our extra selection is effective. Besides, we further outperform the vanilla text-aware method (only text tokens) by $\textbf{2.7}\%$ on POPE (Li et al., 2023b). The huge margin means POPE sparsification is quite sensitive to question prompts, and text guidance is necessary. In summary, text rater selection is general and improves the performance across scenarios.
@@ 5.2 Recycling of Pruned Tokens
To validate the effectiveness of our token recycling strategy, we perform ablation experiments on the LLaVA model (Liu et al., 2024a). The results are presented in Table 4. Across multiple sparsity ratios ($64$, $96$, $128$, $192$), our algorithm achieves a significant average performance improvement of $\textbf{1.2}\%$ and $\textbf{7.2}\%$ on TextVQA (Singh et al., 2019) and POPE (Li et al., 2023b), respectively. Notably, as the number of pruned vision tokens increases, the benefit brought by our recycling method increases. For instance, when pruning from $192$ to $64$ tokens, the pruned token recycling significantly boosts the accuracy from 1.5% to 17.7% on POPE. We argue that when the size of the deleted pool grows, the amount of lost information increases. Our method effectively recycles the lost information and compresses it into few slots using the proposed reconstruction mechanism.
@@ 5.3 Computational Efficiency
SparseVLM affords significant efficiency and storage gains for the inference process. We conduct a comparative analysis of CUDA time, and FLOPs on LLaVA-7B, and compare our method with the baseline method and FastV (Chen et al., 2024a). As displayed in Table 1, we conduct an inference efficiency analysis on a single NVIDIA A100-80GB with identical lengths of text prompts and single-image inputs. Compared to the baseline model, SparseVLM achieves a significant reduction of $43.1\%$ in CUDA time and $62.8\%$ in FLOPs while keeping $96.7\%$ accuracy. Despite SparseVLM has a minimal overhead to calculate text raters and cluster-pruned vision tokens, it leads to fewer than FastV tokens with comparable accuracy. Additionally, SparseVLM saves $67\%$ cache memory compared to vanilla LLaVA (where 302.4MB is reduced to 100.8MB), while keeping $99.1\%$ accuracy. More efficiency visualization (e.g., efficiency on VideoLLaVA) can be found in the Appendix G.